OpenStudy (anonymous):

easy chem q: Given the following unbalanced chemical equation, how many moles of nitrogen monoxide are needed to react with excess ammonia to produce 32.6 grams of water? ___NH3 (g) + ___ NO (g) ___N2 (g) + ___H2O (l) Answer 5.53 x 10-1 moles NO 10.9 moles NO 32.6 moles NO 1.81 moles NO I know the balanced chemical equation is 4 NH3 + 6NO = 5N2 + 6H2O, I can't seem to figure out the rest.

OpenStudy (kainui):

So you've basically worked out that for every 6 mol of NO you use in a reaction you will be getting 6 mol of H2O, right? Another way you could write this is: \[6mol NO=6molH_2O\] This seems weird, but it makes sense, right? If you multiply one side by 2, the other must be multiplied by 2, because for ever mole of NO you make a mole of H2O every time. You can rewrite this like: \[\frac{ 6 mol NO }{ 6molH_2O }=1\] Why would you want to do that? I'll show you in a bit, but remember, you can multiply anything by 1 and it doesn't change it. This is the idea of a conversion factor. Now, on the periodic table you have these numbers that aren't quite whole numbers, for Hydrogen it's about 1.01 and for Oxygen it's about 15.99. Not particularly important to be exact, we can assume H is 1 and O is 16 most of the time just to make the calculation easier on us, but if your professor wants more decimal points of accuracy, it can't hurt. So what do those numbers represent? They represent how many grams 1 mole of that atom weighs. So, oxygen weighs 16 grams per mole. Remember, you are trying to convert from grams of H2O to moles of NO, so hopefully you're seeing the connections here to get you where you need to go. So H is 1 gram per mole, O is 16 g/mol, but you have water which is H2O, so simply add up 2 H and 1 O atom per mole to get 18 grams per mole of water. For consistency, this is true right? \[18 gram H_2O=1molH_2O\] But remember this can still be thought of as:\[\frac{ 18 g H_2O }{ 1molH_2O }=1\] Great, so now you are basically there, just need to put it together: \[32.6gH_2O \frac{ 1mol H_2O }{ 18 g H_2O }\frac{ 6molNO }{ 6molH_2O }=molNO\] See how we start out with what you're given, and then multiply by our conversion factors (just fancy versions of the number 1)? Also, don't just think about dividing the numbers, divide the units too! By that I mean, see how moles of H2O from the first fraction divides out with moles of H2O from the second fraction? It's not that confusing, if you have 2/2=1 but now we're looking at elephants/elephants=1. It doesn't matter that it's not a number, but it helps a lot in making sure everything works out correctly.

OpenStudy (anonymous):

Thank you so much! This makes so much sense now.

OpenStudy (kainui):

Glad I could help, if you have any questions while I'm here, you should try to ask sooner rather than later while this is still on my mind! =D